# A mixed finite element method for a class of fourth-order stochastic evolution equations with multiplicative noise

### v1

@@ 3. A fully-discrete mixed finite element method
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{n}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa)e^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa)e^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa)e^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\left|{I_{9}}\right|+\left|{I_{12}}\right|$ $\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ce^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\leq\epsilon\mathbb{E}\left[\max_{m\leq n}\mathds{1}_{\Omega_{\kappa,m-1}}\left\|{\nabla\boldsymbol{\theta}^{m}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ce^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\leq Ce^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+\epsilon\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}-\nabla\boldsymbol{\theta}^{\ell-1}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
@@
$\leq\mathbb{E}\left[\left\|{\nabla\boldsymbol{\theta}^{0}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa^{4})e^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$ (3.47)
@@
$k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$ $\leq\mathbb{E}\left[\left\|{\nabla\boldsymbol{\theta}^{0}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa^{4})e^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\mathbb{E}\left[\max_{m\leq n}\left(\mathds{1}_{\Omega_{\kappa,m-1}}\left\|{\boldsymbol{u}(t_{m})-\boldsymbol{u}_{h}^{m}}\right\|_{\mathbb{H}^{1}}^{2}\right)\right]+k\sum_{\ell=1}^{n}\mathbb{E}\left[\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{H}(t_{\ell})-\boldsymbol{H}_{h}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right]\leq\widetilde{C}e^{\widetilde{C}\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$ (3.49)
@@
By Chebyshev’s inequality and Theorem 3.7 with $\kappa^{2}=O\left(\log(\log 1/h)\right)$, for any $\alpha,\delta>0$ we have
@@
for any $r\geq 1$. In particular, the right-hand side of (3.52) tends to 00 as $h,k\to 0^{+}$.
@@
$\kappa^{2}=\frac{1}{\widetilde{C}}\left(\left|{\log\left(h^{2}+k^{\frac{1}{2}-\delta}\right)}\right|-(2^{q-1}-1)\log\left|{\log\left(h^{2}+k^{\frac{1}{2}-\delta}\right)}\right|\right),$
@@
$\leq\widetilde{C}e^{\widetilde{C}\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+C_{q}\kappa^{-2^{q-1}}$

### v2

@@ 3. A fully-discrete mixed finite element method
$\quad+\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left|{\left\langle\boldsymbol{u}_{h}^{\ell}\times\left(\Pi_{h}\big(f_{R}(\boldsymbol{u}(t_{\ell}))-f_{R}(\boldsymbol{u}_{h}^{\ell-1})\big)\right),\Delta_{h}\boldsymbol{\theta}^{\ell}\right\rangle}\right|\mathrm{d}s\right]$
$\quad+\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left|{\left\langle\boldsymbol{u}_{h}^{\ell}\times\left(\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})\right),\Delta_{h}\boldsymbol{\theta}^{\ell}\right\rangle}\right|\mathrm{d}s\right]$
$\leq I_{8a}+I_{8b}+I_{8c}+I_{8d}.$
For the terms $I_{8a}$ and $I_{8b}$, we use (2.11) and Young’s inequality to obtain
$\left|{I_{8a}}\right|+\left|{I_{8b}}\right|$ $\leq Ch^{2}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{L^{\infty}_{T}(\mathbb{H}^{3})}^{2}\sum_{\ell=1}^{n}k\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{4}}^{2}\right]$
$\quad+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\leq Ch^{2}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{L^{\infty}_{T}(\mathbb{H}^{3})}^{4}+\left(\sum_{\ell=1}^{n}k\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right)^{2}\right]$
$\quad+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\leq Ch^{2}+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
For $I_{8c}$, we use the stability of $\Pi_{h}$, Lipschitz continuity of $f_{R}$, and Hölder continuity of $\boldsymbol{u}$ to obtain
$\left|{I_{8c}}\right|$ $\leq Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{4}}\left\|{\boldsymbol{u}(t_{\ell})-\boldsymbol{u}(t_{\ell-1})+\boldsymbol{\rho}^{\ell-1}}\right\|_{\mathbb{L}^{2}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}\right]$
$\quad+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}+\mathcal{R}_{h}\boldsymbol{u}(t_{\ell})}\right\|_{\mathbb{L}^{4}}\left\|{\boldsymbol{\theta}^{\ell-1}}\right\|_{\mathbb{L}^{2}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}\right]$
$\leq C(h^{4}+k^{1-\delta})\,\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{1/2-\delta}_{T}(\mathbb{L}^{2})}^{4}+\left(\sum_{\ell=1}^{n}k\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right)^{2}\right]+C\kappa k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right]$
$\quad+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\leq C(1+\kappa)e^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
For $I_{8d}$, by Young’s inequality and Hölder continuity of $\boldsymbol{H}$, we have
$\left|{I_{8d}}\right|$ $\leq\mathbb{E}\left[\sum_{\ell=1}^{n}k\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}\left(k^{\frac{1}{4}-\delta}\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{\alpha}_{T}(\mathbb{H}^{3})}\right)\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}\right]$
$\leq Ck^{\frac{1}{2}-\delta}+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
Altogether, we thus obtain
$\left|{I_{8}}\right|$ $\leq C(1+\kappa)e^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa)e^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa)e^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\left|{I_{9}}\right|+\left|{I_{12}}\right|$ $\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ce^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\leq\epsilon\mathbb{E}\left[\max_{m\leq n}\mathds{1}_{\Omega_{\kappa,m-1}}\left\|{\nabla\boldsymbol{\theta}^{m}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ce^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\leq Ce^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+\epsilon\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}-\nabla\boldsymbol{\theta}^{\ell-1}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
@@
$\leq\mathbb{E}\left[\left\|{\nabla\boldsymbol{\theta}^{0}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ce^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$ (3.47)
@@
$k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$ $\leq\mathbb{E}\left[\left\|{\nabla\boldsymbol{\theta}^{0}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ce^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right).$
@@
$\mathbb{E}\left[\max_{m\leq n}\left(\mathds{1}_{\Omega_{\kappa,m-1}}\left\|{\boldsymbol{u}(t_{m})-\boldsymbol{u}_{h}^{m}}\right\|_{\mathbb{H}^{1}}^{2}\right)\right]+k\sum_{\ell=1}^{n}\mathbb{E}\left[\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{H}(t_{\ell})-\boldsymbol{H}_{h}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right]\leq\widetilde{C}e^{\widetilde{C}\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$ (3.49)
@@
By Chebyshev’s inequality and Theorem 3.7 with $\kappa=O\left(\log(\log 1/h)\right)$, for any $\alpha,\delta>0$ we have
@@
for any $r\geq 1$. In particular, the right-hand side of (3.52) tends to 0 as $h,k\to 0^{+}$.
@@
$\kappa=\frac{1}{\widetilde{C}}\left(\left|{\log\left(h^{2}+k^{\frac{1}{2}-\delta}\right)}\right|-(2^{q-1}-1)\log\left|{\log\left(h^{2}+k^{\frac{1}{2}-\delta}\right)}\right|\right),$
@@
$\leq\widetilde{C}e^{\widetilde{C}\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+C_{q}\kappa^{-2^{q-1}}$
@@
4. Numerical experiments
We present a set of numerical experiments to validate the theoretical convergence properties of the proposed finite element scheme for the sLLBar equation. All computations are carried out in the FEniCS environment. The computational domain $\mathscr{D}$ is taken as the unit square $0,1]\times[0,1]$. The magnetisation vector field $\boldsymbol{u}$ and the effective field $\boldsymbol{H}$ are discretised in space using continuous piecewise linear finite elements on a family of quasi-uniform meshes.
To assess convergence, we compute a reference solution on a very fine mesh and with a small time step. This reference solution serves as an approximation of the exact stochastic solution for each realisation of the Wiener process. For coarser discretisations, we use the same Brownian path so that the difference between the numerical solution and the reference is meaningful pathwise. The errors $\mathcal{E}_{s}^{\boldsymbol{u}}(h,k)$ and $\mathcal{E}_{s}^{\boldsymbol{H}}(h,k)$ at final time $T$ is then measured in $L^{2}(\Omega;\mathbb{H}^{s})$-norm for $s=0$ or $1$, defined by
$\mathcal{E}_{s}^{\boldsymbol{u}}(h,k):=\left(\mathbb{E}\left\|{\boldsymbol{u}_{h}^{N}-\boldsymbol{u}_{\text{ref}}(T)}\right\|_{\mathbb{H}^{s}}^{2}\right)^{\frac{1}{2}}\;\text{ and }\;\mathcal{E}_{s}^{\boldsymbol{H}}(h,k):=\left(\mathbb{E}\left\|{\boldsymbol{H}_{h}^{N}-\boldsymbol{H}_{\text{ref}}(T)}\right\|_{\mathbb{H}^{s}}^{2}\right)^{\frac{1}{2}}.$
Here, $\big(\boldsymbol{u}_{h}^{N},\boldsymbol{H}_{h}^{n}\big)$ is the numerical solution with mesh size $h$ at time $T=Nk$, and $\big(\boldsymbol{u}_{\text{ref}}(T),\boldsymbol{H}_{\text{ref}}(T)\big)$ denotes the reference solution at the same final time along the same realisation of the Brownian motion. In practice, the expectation is approximated by a Monte Carlo average over $M$ independent sample paths.
We vary separately the mesh size $h$ and the time step $k$ to test spatial and temporal convergence. To verify spatial convergence, we fix a sufficiently small time step and compare errors across a sequence of meshes ($h=2^{-j}$, for $j=2,3,4$, with $j=6$ as the reference solution). For temporal convergence, we fix a fine mesh and vary the time step size ($k=(25\times 2^{j})^{-1}$, for $j=0,1,2$, with $j=4$ as the reference solution). In both cases, errors are averaged over $M=25$ realisations. The experimental rate of convergence is obtained by fitting the errors against mesh size or time step in a log-log plot.
4.1. Simulation 1
In this simulation, we take the parameters to be $\lambda_{1}=0.2$, $\lambda_{2}=0.1$, $\gamma=5.0$, $\kappa=0.5$, $\mu=1.0$, $\beta_{1}=0.1$, and $\beta_{2}=0.05$. The current density $\boldsymbol{\nu}=(1,0)^{\top}$. The initial data is specified to be
$\boldsymbol{u}_{0}(x)=\big(-y,x,0\big),$
and the vector field $\boldsymbol{g}$ is taken to be
$\boldsymbol{g}(x)=\big(0.8(1-x),0.2,0.5(1+x)\big).$
We solve the sLLBar equation by employing the implicit scheme (3.1)
Snapshots of a sample path of the magnetisation vector field $\boldsymbol{u}$ and the effective field $\boldsymbol{H}$ with mesh-size $h=1/16$ at selected times are shown in Figures 1 and 3, respectively. The colour indicates the relative value of the magnitude. Figures 2(a) and 2(b) display the plots of $\mathcal{E}_{s}^{\boldsymbol{u}}$ against $1/h$ and $1/k$, respectively. Similar plots for $\mathcal{E}_{s}^{\boldsymbol{H}}$ against $1/h$ and $1/k$ are shown in Figures 4(a) and 4(b). These results are consistent with Theorem 3.7 in the $\mathbb{H}^{1}$ norm, while the numerical simulation indicates an even higher convergence rate in the $\mathbb{L}^{2}$ norm. Such behavior is in line with what is traditionally expected in finite element analysis, though a rigorous proof in our setting remains an interesting open question.
Figure 5 shows the energy of the system over 30 independent sample paths for $h=1/16$, $k=0.01$, and for $h=1/32$, $k=0.005$. We recall that the energy is defined as
$\text{Energy}(\boldsymbol{u}):=\frac{1}{2}\left\|{\nabla\boldsymbol{u}}\right\|_{\mathbb{L}^{2}}^{2}+\frac{\kappa}{4}\left\|{\left|{\boldsymbol{u}}\right|^{2}-\mu}\right\|_{\mathbb{L}^{2}}^{2}.$
For relatively small noise intensity, the energy shows only minor pathwise fluctuations and, on average, decays over time.
(a) $t=0$
(b) $t=0.02$
(c) $t=0.1$
(d) $t=0.2$
Figure 1. Snapshots of a sample path of the magnetisation $\boldsymbol{u}$ in simulation 1.
[…]
(a) Spatial convergence order of $\boldsymbol{u}$.
[…]
(b) Temporal convergence order of $\boldsymbol{u}$.
Figure 2. Convergence orders of magnetisation vector field $\boldsymbol{u}$ in simulation 1.
(a) $t=0$
(b) $t=0.02$
(c) $t=0.1$
(d) $t=0.2$
Figure 3. Snapshots of a sample path of the effective field $\boldsymbol{H}$ in simulation 1.
[…]
(a) Spatial convergence order of $\boldsymbol{H}$.
[…]
(b) Temporal convergence order of $\boldsymbol{H}$.
Figure 4. Convergence orders of effective field $\boldsymbol{H}$ in simulation 1.
(a) Graph of energy vs time with $h=1/16$ and $k=1/100$ for 30 sample paths.
(b) Graph of energy vs time with $h=1/32$ and $k=1/200$ for 30 sample paths.
Figure 5. Energy evolution in simulation 1
4.2. Simulation 2
In this simulation, we take the parameters to be $\lambda_{1}=0.5$, $\lambda_{2}=0.05$, $\gamma=8.0$, $\kappa=0.25$, $\mu=1.0$, $\beta_{1}=0.2$, and $\beta_{2}=0.1$. The current density $\boldsymbol{\nu}=(2,0)^{\top}$. The initial data is specified to be
$\boldsymbol{u}_{0}(x)=\big(\sin(2\pi y),\sin(2\pi x),0\big),$
and the vector field $\boldsymbol{g}$ is taken to be
$\boldsymbol{g}(x)=\big(5(1+x),10(1+y),2\cos(2\pi x)\big).$
Snapshots of a sample path of the magnetisation vector field $\boldsymbol{u}$ and the effective field $\boldsymbol{H}$ with mesh-size $h=1/16$ at selected times are shown in Figures 6 and 8, respectively. The colour indicates the relative value of the magnitude. Figures 7(a) and 7(b) display the plots of $\mathcal{E}_{s}^{\boldsymbol{u}}$ against $1/h$ and $1/k$, respectively. Similar plots for $\mathcal{E}_{s}^{\boldsymbol{H}}$ against $1/h$ and $1/k$ are shown in Figures 9(a) and 9(b).
Figure 10 shows the energy over 30 independent sample paths for $h=1/16$, $k=0.02$, and for $h=1/32$, $k=0.01$. Since the noise intensity is moderately large, the energy trajectory exhibits pronounced pathwise variability around the mean profile.
(a) $t=0$
(b) $t=0.05$
(c) $t=0.1$
(d) $t=0.2$
Figure 6. Snapshots of a sample path of the magnetisation $\boldsymbol{u}$ in simulation 2.
[…]
(a) Spatial convergence order of $\boldsymbol{u}$.
[…]
(b) Temporal convergence order of $\boldsymbol{u}$.
Figure 7. Convergence orders of magnetisation vector field in simulation 2.
(a) $t=0$
(b) $t=0.05$
(c) $t=0.1$
(d) $t=0.2$
Figure 8. Snapshots of a sample path of the effective field $\boldsymbol{H}$ in simulation 2.
[…]
(a) Spatial convergence order of $\boldsymbol{H}$.
[…]
(b) Temporal convergence order of $\boldsymbol{H}$.
Figure 9. Convergence orders of effective field in simulation 2.
(a) Graph of energy vs time with $h=1/16$ and $k=1/50$ for 30 sample paths.
(b) Graph of energy vs time with $h=1/32$ and $k=1/100$ for 30 sample paths.
Figure 10. Energy evolution in simulation 2